OpenStudy (anonymous):

An automobile's radiator has a capacity of fifteen quarts, and it currently contains twelve quarts of a thirty percent antifreeze solution. How many quarts of pure antifreeze must be added to strengthen the solution to forty percent? 2 quarts 3 quarts 4 quarts

hero (hero):

12 + x = y 12(.30) + 1x = y(.40)

hero (hero):

The 15 is a trick

hero (hero):

12 + x = y 3.6 + x = .4y .4(12 + x) = 3.6 + x 4.8 + .4x = 3.6 + x 4.8 - 3.6 = x - .4x 1.2 = .6x 1.2/.6 = x 2 = x

hero (hero):

Therefore, 2 quarts must be added to get the 40%

OpenStudy (anonymous):

thank u